# IMU: Influence-guided Machine Unlearning

### v2

@@ 3.2 IMU
Influence estimation. Bengio, Courville, and Vincent (2013) suggests that high-dimensional data tends to lie on lower-dimensional manifolds. Hence, rather than estimating influence on the full deep network, which would require inverting a high-dimensional, potentially ill-conditioned Hessian, we instead compute influence at the level of learned representations. Specifically, we treat the feature representation ${\bm{z}}=\phi({\bm{x}})$ as the input and estimate the influence of each data point $\bm{x}$} w.r.t. the classifier $h(\cdot)$. Therefore, we estimate the influence value w.r.t. the data point ${\bm{x}}\in{\mathcal{D}}_{f}$ as
${\mathcal{I}}^{c}({\bm{x}})=-{\bm{H}}_{{\bm{\theta}}_{c}^{\ast}}^{-1}\nabla_{{\bm{\theta}}_{c}}\ell\big{(}h_{{\bm{\theta}}_{c}^{\ast}}({\bm{z}})\big{)},$ (3)
where the latent representation ${\bm{z}}=\phi_{{\bm{\theta}}_{e}^{\ast}}({\bm{x}})$. And the change in the loss value for ${\mathcal{D}}_{f}$ when the training data point $\bm{x}$} is removed from the training set can then be approximated by
${\mathcal{I}}^{c}({\bm{x}},{\mathcal{D}}_{f})=-\nabla_{{\bm{\theta}}_{c}}\ell\big{(}h_{{\bm{\theta}}_{c}^{\ast}}(\mathbf{Z}_{f})\big{)}^{\top}{\bm{H}}_{{\bm{\theta}}_{c}^{\ast}}^{-1}\nabla_{{\bm{\theta}}_{c}}\ell\big{(}h_{{\bm{\theta}}_{c}^{\ast}}({\bm{z}})\big{)},$ (4)
where $\mathbf{Z}_{f}=\phi_{{\bm{\theta}}_{e}^{\ast}}({\mathcal{D}}_{f})$. This can significantly improve the stability and accuracy of influence estimation and avoid the computational cost of full Hessian inversion. Figure 1 presents examples with different levels of influence on the data from their respective class. Top influences are examples whose omission from the training data is predicted to most increase the loss on the corresponding samples. In Figure 2, we show that equally treating these examples (denoted as GA) in unlearning potentially causes either under-unlearning or over-unlearning. Hence, in the following, we introduce the influence-guided loss to dynamically allocate attention for different data points.
@@
3:  Compute ${\bm{H}}_{{\bm{\theta}}_{c}}=\mathbb{E}_{{\mathcal{D}}_{f}}\left[\nabla^{2}_{{\bm{\theta}}_{c}}\ell\left(h_{{\bm{\theta}}_{c}}\left({\bm{z}}_{i}\right)\right)\right]$ and ${\bm{Z}}_{f}=\phi_{{\bm{\theta}}_{e}}({\mathcal{D}}_{f})$ where ${\bm{z}}_{i}=\phi_{{\bm{\theta}}_{e}}({\bm{x}}_{i}),\forall{\bm{x}}_{i}\sim{\mathcal{D}}_{f}$.
4:  $\forall{\bm{x}}\sim{\mathcal{D}}_{f}$, estimate influence value ${\mathcal{I}}^{c}({\bm{x}},{\mathcal{D}}_{f})$ via Eq. 4: $\nabla_{{\bm{\theta}}_{c}}\ell\big{(}h_{{\bm{\theta}}_{c}}(\mathbf{Z_{f}})\big{)}^{\top}{\bm{H}}_{{\bm{\theta}}_{c}}^{-1}\nabla_{{\bm{\theta}}_{c}}\ell\big{(}h_{{\bm{\theta}}_{c}}({\bm{z}})\big{)}$.
5:  Select the samples satisfying ${\mathcal{I}}^{c}({\bm{x}}_{i},{\mathcal{D}}_{f})<0$ and process unlearning with these samples ${\mathcal{D}}_{f}^{{\mathcal{I}}^{c}<0}$.
6:  Compute the influence-guided loss ${\mathcal{L}}_{\texttt{IMU}}({\mathcal{D}}_{f};{\bm{\theta}}^{(t)})$ via Eq. 5: $\mathbb{E}_{{\bm{x}}_{i}\sim{\mathcal{D}}_{f}^{{\mathcal{I}}^{c}<0}}\left[\hat{{\mathcal{I}}^{c}}({\bm{x}}_{i},{\mathcal{D}}_{f})\cdot\ell\left(f_{{\bm{\theta}}}\left({\bm{x}}_{i}\right)\right)\right]$.
@@
Influence-guided MU. To realize dynamically allocating different degrees of attention to data points according to their influence values, we integrate their corresponding influence values into the MU process by defining an influence-guided loss as
${\mathcal{L}}_{\texttt{IMU}}({\mathcal{D}}_{f};{\bm{\theta}})$ $=-\mathbb{E}_{{\bm{x}}_{i}\sim{\mathcal{D}}_{f}}\left[\hat{{\mathcal{I}}^{c}}({\bm{x}}_{i},{\mathcal{D}}_{f})\cdot\ell\big{(}f_{\bm{\theta}}({\bm{x}}_{i})\big{)}\right],$ (5)
where $\hat{{\mathcal{I}}^{c}}({\bm{x}}_{i},{\mathcal{D}}_{f})=\left|\sqrt{{\mathcal{I}}^{c}({\bm{x}}_{i},{\mathcal{D}}_{f})}\right|/\sum_{j}\left|\sqrt{{\mathcal{I}}^{c}({\bm{x}}_{j},{\mathcal{D}}_{f})}\right|$ denotes the normalized score. Different from the naive gradient ascent, Equation 5 is equivalent to taking a weighted average of the per-sample losses, with weights proportional depending on their influence function value. Whereas, in practice, due to the large variance of the influence value distribution of different samples and the error of approximate estimation, the influence value of a certain sample can be very large. Such extreme value can dominate the weighted average, effectively diminishing the relative contribution of others that are equally important, leading the model to overlook them. To tackle this problem, we try to smooth and truncate after computing ${\bm{I}}=\begin{bmatrix}{\mathcal{I}}^{c}({\bm{x}}_{1},{\mathcal{D}}_{f}),\cdots,{\mathcal{I}}^{c}({\bm{x}}_{m},{\mathcal{D}}_{f})\end{bmatrix}^{\top}$. If ${\mathcal{I}}^{c}({\bm{x}}_{i},{\mathcal{D}}_{f})<0$, it indicates that this sample is helpful in improving the model’s performance on ${\mathcal{D}}_{f}$. Therefore, we only need to select these samples with a negative impact to participate in unlearning. Algorithm 1 describes the procedure of our algorithm IMU in detail. We use these influence scores to dynamically adjust the unlearning strategy, assigning sample-specific attention during the parameter update process. Samples with higher influence values receive more aggressive updates, enabling targeted and efficient forgetting while preserving the performance on retain and unseen data.
Figure 2: Comparison between GA, NPO, and IMU.
3.3 Connection with NPO
Negative Preference Optimization (NPO) (Zhang et al. 2024) stands out in LLM unlearning, it defines
${\mathcal{L}}_{\texttt{NPO}}({\bm{\theta}})\coloneqq\frac{2}{\beta}\mathbb{E}_{{\mathcal{D}}_{f}}\left[\log\left(1+\left(\frac{\pi_{{\bm{\theta}}}({\bm{y}}|{\bm{x}})}{\pi_{\texttt{ref}}({\bm{y}}|{\bm{x}})}\right)^{\beta}\right)\right],$ (6)
where $\pi_{{\bm{\theta}}}({\bm{y}}|{\bm{x}})$ denotes the output probability of the current model, and $\pi_{\texttt{ref}}({\bm{y}}|{\bm{x}})$ is the output probability from a reference model (e.g., the original model). For image classification, the model typically outputs a softmax probability distribution over classes as $\pi_{{\bm{\theta}}}({\bm{y}}|{\bm{x}})=\texttt{softmax}\left(f_{{\bm{\theta}}}({\bm{x}})\right)$, where $f_{{\bm{\theta}}}({\bm{x}})$ are the logits, and $\ell({\bm{x}},{\bm{y}};{\bm{\theta}})=-\log\pi_{{\bm{\theta}}}({\bm{y}}|{\bm{x}})$.
Refer to Proposition 1 in NPO (Zhang et al. 2024), NPO is a strict generalization of the gradient ascent (GA) method. The gradients of GA, NPO, and IMU are
$\nabla_{{\bm{\theta}}}{\mathcal{L}}_{\texttt{GA}}=-\mathbb{E}_{({\bm{x}},{\bm{y}})\sim{\mathcal{D}}_{f}}\left[\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{y}};{\bm{\theta}})\right],$ (7)
$\nabla_{{\bm{\theta}}}{\mathcal{L}}_{\texttt{NPO}}=-\mathbb{E}_{({\bm{x}},{\bm{y}})\sim{\mathcal{D}}_{f}}\left[{\mathcal{W}}_{{\bm{\theta}}}({\bm{x}},{\bm{y}})\cdot\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{y}};{\bm{\theta}})\right],$ (8)
$\nabla_{{\bm{\theta}}}{\mathcal{L}}_{\texttt{IMU}}=-\mathbb{E}_{({\bm{x}},{\bm{y}})\sim{\mathcal{D}}_{f}}\left[\hat{{\mathcal{I}}^{c}}({\bm{x}},{\mathcal{D}}_{f})\cdot\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{y}};{\bm{\theta}})\right],$ (9)
where ${\mathcal{W}}_{{\bm{\theta}}}({\bm{x}},{\bm{y}})=2\pi_{{\bm{\theta}}}({\bm{y}}|{\bm{x}})^{\beta}/\left(\pi_{{\bm{\theta}}}({\bm{y}}|{\bm{x}})^{\beta}+\pi_{\texttt{ref}}({\bm{y}}|{\bm{x}})^{\beta}\right)$ is the adaptive smoothing weight, controlling the divergence speed in unlearning. NPO adaptively assigns weights to data points to help prevent catastrophic collapse in GA; however, Fan et al. (2024a) shows that NPO exhibits reference model bias, causing an uneven allocation of unlearning power. IMU instead uses explicitly estimated influence scores to guide the unlearning process. In this sense, IMU can be viewed as an influence-weighted counterpart to GA, offering a more direct and interpretable mechanism for prioritizing which data points to forget. Figure 2 presents the class-wise forgetting on CIFAR-10. We can observe that the forget quality of NPO decreases to 2.37% quickly and then stabilizes in later epochs, but it is accompanied by a continuous and rapid decline in model utility, from 92.01% to 85.89%. In contrast, the forget accuracy of IMU has decreased more slowly and has consistently maintained a high test accuracy around 92.60%. Further comparisons in language tasks can be found in § 4, and divergence analysis can be found in the appendix.
4 Experiment
We empirically evaluate IMU to answer the following research questions (RQs): RQ1 How well can IMU balance forget quality and model utility across various unlearning scenarios? RQ2 How well does IMU generalize across different tasks and datasets? RQ3 How efficient is IMU?
4.1 Setup

### v3

@@ 3.2. IMU
Influence estimation. According to the manifold hypothesis, high-dimensional data tends to lie on lower-dimensional manifolds (Bengio et al., 2013). Hence, rather than estimating influence on the full deep network, which would require inverting a high-dimensional, potentially ill-conditioned Hessian, we instead compute influence at the level of learned representations. Specifically, we treat the feature representation ${\bm{z}}=\phi({\bm{x}})$ as the input and estimate the influence of $\bm{x}$} w.r.t. the classifier $h(\cdot)$. We’ll formally justify this approximation in § 3.3.
Removing $\bm{x}$} corresponds to setting $\epsilon=-1/N$, therefore, we estimate the parameter change under removal as
(3) $\Delta{\bm{\theta}}^{\ast}_{c,-{\bm{x}}}\approx-\frac{1}{N}{\mathcal{I}}^{c}_{\texttt{up},{\bm{\theta}}^{\ast}_{c}}({\bm{x}})=\frac{1}{N}{\bm{H}}_{{\bm{\theta}}_{c}^{\ast}}^{-1}\nabla_{{\bm{\theta}}_{c}}\ell\big(h_{{\bm{\theta}}_{c}^{\ast}}({\bm{z}})\big),$
where the latent representation ${\bm{z}}=\phi_{{\bm{\theta}}_{e}^{\ast}}({\bm{x}})$. Accordingly, the change in the loss over ${\mathcal{D}}_{f}$ induced by removing $\bm{x}$} is
$\Delta_{-{\bm{x}}}$ $\ell({\mathcal{D}}_{f})\approx(-1/N)\;{\mathcal{I}}^{c}_{\texttt{up},\ell}({\bm{x}},{\mathcal{D}}_{f})$
(4) $=$ $(1/N)\;\nabla_{{\bm{\theta}}_{c}}\ell\big(h_{{\bm{\theta}}_{c}^{\ast}}(\mathbf{Z}_{f})\big)^{\top}{\bm{H}}_{{\bm{\theta}}_{c}^{\ast}}^{-1}\nabla_{{\bm{\theta}}_{c}}\ell\big(h_{{\bm{\theta}}_{c}^{\ast}}({\bm{z}})\big),$
where $\mathbf{Z}_{f}=\phi_{{\bm{\theta}}_{e}^{\ast}}({\mathcal{D}}_{f})$. Then $\Delta_{-{\bm{x}}}\ell({\mathcal{D}}_{f})>0$ indicates that $\bm{x}$} is helpful for learning ${\mathcal{D}}_{f}$, whereas $\Delta_{-{\bm{x}}}\ell({\mathcal{D}}_{f})<0$ indicates that $\bm{x}$} is harmful to ${\mathcal{D}}_{f}$.
Given the access constraint (${\mathcal{D}}_{r}$ is unavailable), computing the exact full-data Hessian is infeasible. Therefore, we use ${\bm{H}}^{f}_{{\bm{\theta}}_{c}^{\ast}}\triangleq\nabla^{2}_{{\bm{\theta}}_{c}}\ell(h_{{\bm{\theta}}_{c}^{\ast}}(\mathbf{Z}_{f}))$, which captures the local curvature of the forgetting objective around ${\bm{\theta}}_{c}^{\ast}$. While this omits the global curvature contributed by ${\mathcal{D}}_{r}$, it explicitly captures the local geometric structure of the forgetting subspace. This yields a targeted influence estimate that is sufficient for our purpose of weighting ${\bm{x}}\sim{\mathcal{D}}_{f}$.
Influence-guided MU. To dynamically allocate unlearning budgets based on sample influence, we integrate these influence scores into the MU process and define the following influence-guided unlearning loss:
(5) ${\mathcal{L}}_{\texttt{IMU}}({\mathcal{D}}_{f};{\bm{\theta}})$ $=-\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{f}^{+}}\hat{w}_{i}\cdot\ell\big(f_{\bm{\theta}}({\bm{x}}_{i})\big),$
where $\hat{w}_{i}$ is the normalized influence scores for sample ${\bm{x}}_{i}$ and is computed as:
(6) $\hat{w}_{i}=\frac{\psi(w_{i})}{\sum_{{\bm{x}}_{j}\in{\mathcal{D}}_{f}^{+}}\psi(w_{j})},w_{i}=\max\!\big(0,\,\Delta_{-{\bm{x}}_{i}}\ell({\mathcal{D}}_{f})\big).$
Since our objective is to remove knowledge about ${\mathcal{D}}_{f}$, we prioritize the helpful data points ($\Delta_{-{\bm{x}}}\ell({\mathcal{D}}_{f})>0$) whose removal would naturally increase the loss and restrict the weighting to the subset ${\mathcal{D}}_{f}^{+}\triangleq\{{\bm{x}}_{i}\in{\mathcal{D}}_{f}\mid\Delta_{-{\bm{x}}}\ell({\mathcal{D}}_{f})>0\}$.
@@
3:  Compute ${\bm{H}}_{{\bm{\theta}}_{c}}=\mathbb{E}_{{\bm{z}}_{i}\in{\mathcal{D}}_{f}}\left[\nabla^{2}_{{\bm{\theta}}_{c}}\ell\left(h_{{\bm{\theta}}_{c}}\left({\bm{z}}_{i}\right)\right)\right]$ where ${\bm{z}}_{i}=\phi_{{\bm{\theta}}_{e}}({\bm{x}}_{i}),\forall{\bm{x}}_{i}\sim{\mathcal{D}}_{f}$. /* Estimate the parameter change under removal */
4:  Estimate $\Delta_{-{\bm{x}}}\ell({\mathcal{D}}_{f})$:   $(1/N)\;\nabla_{{\bm{\theta}}_{c}}\ell\big(h_{{\bm{\theta}}_{c}}(\mathbf{Z}_{f})\big)^{\top}{\bm{H}}_{{\bm{\theta}}_{c}}^{-1}\nabla_{{\bm{\theta}}_{c}}\ell\big(h_{{\bm{\theta}}_{c}}({\bm{z}})\big)$. /* Process with samples that are helpful to ${\mathcal{D}}_{f}$ */
5:  Set $w_{i}=\max\!\big(0,\,\Delta_{-{\bm{x}}_{i}}\ell({\mathcal{D}}_{f})\big),\hat{w}_{i}=\frac{\psi(w_{i})}{\sum_{\in{\mathcal{D}}_{f}^{+}}\psi(w_{j})},$ where ${\mathcal{D}}_{f}^{+}\triangleq\{{\bm{x}}_{i}\in{\mathcal{D}}_{f}\mid\Delta_{-{\bm{x}}}\ell({\mathcal{D}}_{f})>0\}$. /* Dynamically allocates unlearning strength */
6:  Compute the influence-guided loss ${\mathcal{L}}_{\texttt{IMU}}({\mathcal{D}}_{f};{\bm{\theta}}^{(t)})$:   $\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{f}^{+}}\hat{w_{i}}\cdot\ell\left(f_{{\bm{\theta}}}\left({\bm{x}}_{i}\right)\right)$.
@@
The overall procedure of IMU is summarized in Algorithm 1. IMU uses these estimated influence scores to dynamically allocate unlearning budgets across the forgetting set. Samples with higher influence scores, which indicate a stronger contribution to the memorization of ${\mathcal{D}}_{f}$, receive proportionally more aggressive updates. This sample-specific attention enables targeted and efficient forgetting while preserving the model’s performance on retain and unseen data. To theoretically ground our approach, we dedicate the following subsection to formally justifying our core approximation: estimating the influence at the classifier level.
3.3. Validity of Classifier-level Approximation
Theoretical analysis. A key design choice in IMU is to estimate influence exclusively at the classifier layer. This choice is motivated by both practical and statistical considerations: full-model influence estimation in DNNs is computationally expensive due to high-dimensional Hessians, and is often unstable due to the non-convexity. We therefore ask whether classifier-level estimation can serve as a reliable surrogate. Motivated by extensive empirical findings in transfer learning, lower-layer representations in well-converged networks tend to be general-purpose and stable, while task-specific adaptation is primarily concentrated in higher layers or the classifier head (Sagun et al., 2018; Yosinski et al., 2014; Donahue et al., 2013; Howard and Ruder, 2018; Kornblith et al., 2019). Building on these established phenomena, we formulate 1 as a local analytical abstraction to theoretically justify our approach.
Assumption 1 (Feature Stability and Weak Coupling).
Let the model parameters be partitioned as ${\bm{\theta}}=[{\bm{\theta}}_{e}^{\top},{\bm{\theta}}_{c}^{\top}]^{\top}$, where ${\bm{\theta}}_{e}$ and ${\bm{\theta}}_{c}$ denote the parameters of the feature extractor and the classifier, respectively. Let ${\mathcal{L}}_{f}({\bm{\theta}})\triangleq\ell(f_{\bm{\theta}}({\mathcal{D}}_{f}))$, and let its Hessian at the converged parameter $\bm{\theta}$$^{*}$ be block-partitioned accordingly as
(7) ${\bm{H}}^{f}_{{\bm{\theta}}^{*}}=\begin{bmatrix}{\bm{H}}_{ee}&{\bm{H}}_{ec}\\ {\bm{H}}_{ce}&{\bm{H}}_{cc}\end{bmatrix}$,
Assuming the presence of convex regularization or Hessian damping standard in influence function estimation (Koh and Liang, 2017), we assume that:
  1. (1)
    ${\bm{H}}_{ee}$ and ${\bm{H}}_{cc}$ are strictly positive definite and invertible, such that $\|{\bm{H}}_{ee}^{-1}\|$ and $\|{\bm{H}}_{cc}^{-1}\|$ are ${\mathcal{O}}(1)$;
  2. (2)
    the feature-classifier Hessian coupling is weak, satisfying $\|{\bm{H}}_{ce}{\bm{H}}_{ee}^{-1}{\bm{H}}_{ec}\|\ll 1;$
  3. (3)
    the feature-level gradients of both the forgetting objective and an individual sample are sufficiently small,
    i.e., $\|\nabla_{{\bm{\theta}}_{e}}{\mathcal{L}}_{f}({\bm{\theta}}^{*})\ll 1\|$ and $\|\nabla_{{\bm{\theta}}_{e}}\ell(f_{{\bm{\theta}}^{*}}({\bm{x}}))\|\ll 1.$
Remark 3.1.
The aforementioned conditions are standard analytical abstractions in both second-order optimization and influence function literature. Specifically, the strictly positive definite condition via Hessian damping (Condition 1) is a ubiquitous prerequisite for ensuring tractable influence estimation in deep networks (Koh and Liang, 2017). Furthermore, the weak feature-classifier coupling (Condition 2) formally encapsulates the block-diagonal Hessian approximations fundamentally relied upon by scalable second-order methods, such as K-FAC (Martens and Grosse, 2015; Donahue et al., 2013). Finally, the small gradient assumption (Condition 3) naturally follows from the standard first-order optimality condition near a converged local minimum (Nocedal and Wright, 2006; Kornblith et al., 2019).
Under 1, we can establish the following theorem, which directly matches the removal score used in Section 3.2. The proof is deferred to § A.1 in the appendix.
Theorem 3.2 (Approximation Error).
Denote the full-model and classifier-level removal scores for a sample $\bm{x}$} as $\Delta^{\mathrm{full}}_{-{\bm{x}}}\ell({\mathcal{D}}_{f})$ and $\Delta^{c}_{-{\bm{x}}}\ell({\mathcal{D}}_{f})$, which is defined respectively as:
(8) $\Delta^{\mathrm{full}}_{-{\bm{x}}}\ell({\mathcal{D}}_{f})$ $\triangleq\frac{1}{N}\nabla_{{\bm{\theta}}}{\mathcal{L}}_{f}({\bm{\theta}}^{*})^{\top}({\bm{H}}^{f}_{{\bm{\theta}}^{*}})^{-1}\nabla_{{\bm{\theta}}}\ell(f_{{\bm{\theta}}^{*}}({\bm{x}})),$
(9) $\Delta^{c}_{-{\bm{x}}}\ell({\mathcal{D}}_{f})$ $\triangleq\frac{1}{N}\nabla_{{\bm{\theta}}_{c}}{\mathcal{L}}_{f}({\bm{\theta}}^{*})^{\top}{\bm{H}}_{cc}^{-1}\nabla_{{\bm{\theta}}_{c}}\ell(f_{{\bm{\theta}}^{*}}({\bm{x}})).$
Under 1, the absolute approximation error is bounded by:
$\Big|\Delta^{\mathrm{full}}_{-{\bm{x}}}\ell({\mathcal{D}}_{f})-\Delta^{c}_{-{\bm{x}}}\ell({\mathcal{D}}_{f})\Big|$
$=\mathcal{O}\!\Bigg($ $\frac{1}{N}\Big(\|\nabla_{{\bm{\theta}}_{e}}{\mathcal{L}}_{f}({\bm{\theta}}^{*})\|\,\|{\bm{H}}_{ee}^{-1}{\bm{H}}_{ec}\|+\|{\bm{H}}_{ce}{\bm{H}}_{ee}^{-1}\|\,\|\nabla_{{\bm{\theta}}_{e}}\ell(f_{{\bm{\theta}}^{*}}({\bm{x}}))\|$
(10) $+\|$ $\nabla_{{\bm{\theta}}_{e}}{\mathcal{L}}_{f}({\bm{\theta}}^{*})\|\,\|\nabla_{{\bm{\theta}}_{e}}\ell(f_{{\bm{\theta}}^{*}}({\bm{x}}))\|+\|{\bm{H}}_{ce}{\bm{H}}_{ee}^{-1}{\bm{H}}_{ec}\|\Big)\Bigg)\ll 1.$
Note that every component constituting the right-hand side of Theorem 3.2 is small under 1, consequently, the approximation gap satisfies $|\Delta^{\mathrm{full}}_{-{\bm{x}}}\ell({\mathcal{D}}_{f})-\Delta^{c}_{-{\bm{x}}}\ell({\mathcal{D}}_{f})|\ll 1$, analytically certifying that the classifier-level score in Section 3.2 faithfully preserves the dominant signal of the full-model removal score.
Figure 2. Relative discrepancy between our computationally efficient classifier-level estimation and the exact full-model influence across individual samples.
Empirical validation. We further empirically evaluate the classifier-level approximation. Because IMU utilizes influence scores predominantly for sample ranking rather than exact magnitude estimation, the critical criterion is whether the computationally efficient classifier-level score preserves the relative ordering induced by the full-model influence. As visually depicted in Figure 2, the two estimations exhibit alignment across samples. Furthermore, when evaluated across five independent random seeds, the classifier-level approximation achieves a striking mean Spearman’s rank correlation coefficient of $0.992$ and a Kendall’s coefficient of concordance of $0.984$. These rank-correlation metrics empirically validate that our localized approximation is not only theoretically sound but also highly effective and robust to stochastic training variations.
Remark 3.3.